A room with 4 people in it all have a birthday in March, which has 31 days in it. What ia the probability that at least two people have their birthday on the same day in March? Assume birthdays are uniformly distributed.

Hint: What is the probability that they all have different birthdays? How many permutations are there when selecting 4 different days out of 31? How many different sets of birthdays are there for these 4 people?
Answer #1

probability that at least two people have their birthday on the same day in March

=1-P(all 4 have birthday on different days)=1-(31*29*28*27)/314=0.182303